please help me solve this word problem please thank you
With only the small valve open, all of the liquid can be drained from a large vat in 4 hours. With only a large valve open, all of the liquid can be drained from the same vat in 2 hours. How long would it take to drain the vat with both valves open?

Answer by Alan3354(31509)

please help me solve this word problem please thank you
With only the small valve open, all of the liquid can be drained from a large vat in 4 hours. With only a large valve open, all of the liquid can be drained from the same vat in 2 hours. How long would it take to drain the vat with both valves open?
----------------------
This type of problem comes up often.
You have to find a common measurement to add the 2 components. Adding the 2 and 4 for 6 doesn't work. It's obvious it will be less than 2 hours.2
-------------------
The "trick" is to invert the times per vat to get work per hour. One valve does 1/2 of the vat per hour, the other does 1/4 per hour.
Add 1/2 adn 1/4 = 3/4 vat per hour together.
Then the inverse of that is 4/3 hours per vat, = 1 hour 20 minutes.
------------------
A shortcut is product over sum:
2*4/(2+4) = 8/6 = 4/3 hours.